OpenStudy (anonymous):

How do I solve 8y - 3(4 - 2y)=6(y+1) Is the answer 144?

OpenStudy (campbell_st):

distribute 1st 8y - 12 + 6y = 6y + 6 then collect like terms

OpenStudy (campbell_st):

and I don't think it's 144...

OpenStudy (yb1996):

First you distribute everything. This way you will get the following equation: 8y - 12 + 6y = 6y +6 Ten you put all of the like terms on one side of the equation. This way you will get: 8y = 18 Then you divide both sides by 8, and you get: y = 18/8 Now you simplify your answer: y = 9/4